namelesss:

A table costs $50 more than a chair. The cost of 6 chairs and 1 table is $750. The equation 6x + x + 50 = 750, where x is the cost of one chair, represents this situation. Plug in the values from the set {50, 100, 150} to find the correct value of x. The value of x that makes the equation true is . So, the cost of a chair is and the cost of a table is .

CripQUEZZ:

Why is this in music?

MxxnLight:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @CripQUEZZ Why is this in music? \(\color{#0cbb34}{\text{End of Quote}}\) right- ;--;

CripQUEZZ:

Well anyways first subtract 50 from both sides 50 - 50 = 750 - 50 The 50 cancels out and your left with 6x + x = 700 Now add your common variables 6x + x = 7x Now your left with 7x = 700 Now just do the opposite of what 7x is doing So you divide 7 from each side 7x / 7 = 700/7 So the seven is canceled out and now your left with x = 700/7 So then simplify 700/7 10 So that means x = 10

CripQUEZZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @CripQUEZZ Well anyways first subtract 50 from both sides 50 - 50 = 750 - 50 The 50 cancels out and your left with 6x + x = 700 Now add your common variables 6x + x = 7x Now your left with 7x = 700 Now just do the opposite of what 7x is doing So you divide 7 from each side 7x / 7 = 700/7 So the seven is canceled out and now your left with x = 700/7 So then simplify 700/7 10 So that means x = 10 \(\color{#0cbb34}{\text{End of Quote}}\) I mean X = 100

CripQUEZZ:

Now jus solve how much they each cost